OpenStudy (anonymous):

Write the equation in standard form and then in slope intercept form: (3, 4) and (5, 8)

OpenStudy (tkhunny):

Slope First. Go!

OpenStudy (anonymous):

I'm not sure where to start since the other problems in this chapter didn't have 2 different points.

OpenStudy (tkhunny):

Have you worked with the SLOPE? Change in y / Change in x?

OpenStudy (anonymous):

Oh yeah, I see. It would be m= 8-4/5-3 = 4/2. But how does the equation end up being: 2x-y=2 in the answer key?

OpenStudy (tkhunny):

Careful. You mean (8-4)/(5-3). The parentheses are NOT optional. The answer key is of no concern. Why are you even looking at that? Let's just keep working before we get jumpy. We have a slope of 2. Now we can write the equation in Point-Slope form. Do you have this form?

OpenStudy (anonymous):

In point slope form I think it looks like this: 8=5x+b

OpenStudy (anonymous):

y=mx+b and y-y1=m(x-x1)

OpenStudy (tkhunny):

y = mx + b is Slope-Intercept. That will do if that's what you have. The second is Slope-Intercept. Substitute into that one.

OpenStudy (anonymous):

point slope form is the second equation I wrote

OpenStudy (tkhunny):

Right. You have a slope, m = 2. You have a point, (3,4) or (5,8). Just pick one and substitute into the Point-Slope Form.

OpenStudy (tkhunny):

* Yea, not sure why I said "Slope-Interecpt" twice. Good call.

OpenStudy (anonymous):

point slope form 8-4=m(5-3) 4=m(2) m=2 y=2x+b 4=2(3)+b 4=6+b b=-2 y=2x-2

OpenStudy (anonymous):

y=2x-2 is standard form

OpenStudy (anonymous):

you can rewrite that as 2x-y=2

OpenStudy (anonymous):

I just don't see how it becomes 2x-y=2. When I wrote it in standard form, it came out as: y=2x+4 (I used the 4/2).

OpenStudy (anonymous):

its supposed to be y=2x-2 you need to use point intercept form and sub in the points given to find the slope and then sub in the slope in standard form and find the y-intercept. Once you have y-intercept you can sub that in and you have the equation then just rearrange it.

OpenStudy (anonymous):

Ok I understand now, thank you!

OpenStudy (anonymous):

your welcome! :)